# Continuous cloud position spectroscopy using a magneto-optical trap

### v1

@@ Abstract
We demonstrate a continuous spectroscopy technique with frequency sensitivity well below the natural transition linewidth, while maintaining a locking range hundreds of times larger. The method exploits the position dependence of a continuous, broadband magneto-optical trap operating on the 7.5 kHz-wide intercombination line of strontium. We show that the frequency sensitivity is fundamentally insensitive to the effective MOT laser linewidth. By applying active feedback on the MOT position to a dispersion-optimized frequency comb, which serves as the reference for stabilizing the MOT laser [1], we achieve a frequency instability below $4.4\times 10^{-13}$ after 400 s of averaging in both the optical and radio-frequency domains, surpassing the performance of conventional hot-vapor modulation transfer spectroscopy. Our method is a broadly applicable alternative route to long-term frequency references in the high $10^{-14}$ range around 100 s.
@@
Continuously operating atomic clocks, such as hydrogen masers in the RF domain [2, 3], and systems employing modulation transfer spectroscopy in the optical domain [4, 5, 6, 7], have demonstrated frequency stability on the order of $\sim 10^{-14}$ after 100 s of averaging [4, 5, 6, 7, 8]. Recent advances have enabled the development of compact, field-deployable optical atomic clocks, including systems based on tetrahedral and pyramidal magneto-optical traps (MOTs) [9, 10, 11], microcell-based clocks [12, 13], and vapor-based clocks [14, 15]. While modulation transfer spectroscopy remains widely used in these systems, it is limited by short-term probe laser instability and a narrow frequency locking range, both typically on the order of the transition linewidth [16, 17]. These limitations are especially pronounced for narrow-line optical transitions, such as those found in the alkaline-earth and alkaline-earth-like atoms. Additionally, modulation transfer spectroscopy does not provide direct access to the RF domain.
@@
In this Letter, we present a simple, continuous spectroscopy technique that exploits the dependence of the MOT position along the vertical axis (gravity) on the MOT laser frequency in a five-beam configuration. We demonstrate that this method improves both the frequency locking range and tolerance to short-term probe laser instability by two orders of magnitude compared to the natural linewidth, while simultaneously achieving a spectral resolution more than an order of magnitude below the 7.5 kHz natural linewidth of the strontium intercombination transition. By applying active feedback based on the MOT position to a dispersion-optimized frequency comb, used for stabilizing the MOT laser, we extract both an 800 MHz RF reference and an optical reference at 689 nm, reducing frequency instabilities below $4.4\times 10^{-13}$ after 400 s of averaging. Finally, we demonstrate that this technique can outperform modulation transfer spectroscopy with hot atomic vapors in terms of long-term frequency stability on the same transition.
@@
Theory. To model the vertical dynamics of the five-beam MOT, where in comparison to a standard MOT the sixth beam propagating in gravity direction is removed [1], we consider the balance between gravitational force $mg$ and optical force along the vertical axis for a single atom with mass $m$ [18]. The resulting steady-state condition determines the MOT position as a function of laser detuning and magnetic field gradient [19, 20]. For strontium, using the $\rm{}^{1}S_{0}\rightarrow\,^{3}P_{1}$ intercombination line at 689 nm with a narrow linewidth of $\gamma_{0}=2\pi\cdot 7.5\,\mathrm{kHz}$, power broadening alone is insufficient to achieve large capture velocities. Instead, frequency modulation is used to extend the effective capture range by more than an order of magnitude. While this “broadband (BB) MOT” technique increases the capture velocities, it cannot be modeled simply as a power-broadened single-frequency (SF) light force. In our implementation, the RF drive of an acousto-optic modulator (AOM) is modulated from $\delta_{\text{start}}$ to $\delta_{\text{start}}-2\pi f_{\text{bandwidth}}$, where $f_{\text{bandwidth}}$ is the modulation bandwidth. In the low-saturation regime (see Appendix A.4), the total force in the vertical direction is:
@@
where $n=f_{\text{bandwidth}}/f_{\text{mod}}$ where $f_{\text{mod}}$ is the modulation frequency, $\sum_{i=0}^{n}\alpha_{i}=1$ are the relative intensities of the $i^{\text{th}}$ frequency component and $I_{0}$ the saturation intensity. For $m_{J}=\pm 1$ of this transition, the Zeeman shift is $\frac{\mu_{B}g_{J}}{h}=2101~\text{kHz/G}$ [21], where $\mu_{B}$ is the Bohr magneton and $g_{J}$ is the Landé g-factor. In equilibrium, $F_{\text{total}}=0$, and the MOT position is directly related to the Zeeman shift. Since we operate close to the center of a quadrupole magnetic field with a linear gradient $B^{\prime}$, a change $\delta_{\text{start}}\rightarrow\delta_{\text{start}}+2\pi\Delta f$ results in a corresponding vertical displacement $\Delta z$, where the frequency-to-position conversion factor $\Delta f/\Delta z=\frac{\mu_{B}g_{J}B^{\prime}}{\hbar}$ is independent of natural linewidth $\gamma_{0}$ and power broadening.
@@
Figure 1: Experimental schematic: (a) Atoms (green disks) are continuously loaded from an upper 2D blue MOT ((b), operating on the $\rm{}^{1}S_{0}\rightarrow^{1}P_{1}$ transition at 461 nm) into a five-beam broadband red MOT ((b), operating on the narrow $\rm{}^{1}S_{0}\rightarrow^{3}P_{1}$ transition at 689 nm). The red MOT is monitored non-destructively via fluorescence imaging, and the center position along the vertical (gravity) axis is extracted every 50 ms. The red MOT laser is locked to an optical frequency comb, which provides short-term frequency stability, while long-term stability is achieved via the MOT position where the vertical displacement $\delta z$ serves as an error signal. Feedback is applied through a tunable RF oscillator used to synchronize the comb, which in turn transfers its stability to both the RF and optical domains.
@@
Experimental Setup. We employ our previously developed architecture for generating a narrow-line strontium MOT using a dispersion-optimized frequency comb-locked laser system, without the need for a free space high-finesse optical cavity [1]. In contrast to typical experiments that execute cooling stages time-sequentially, we execute them in a spatial sequence. Atoms effuse from an oven, are Zeeman slowed and then cooled in a blue 2D MOT, out of which they continuously fall downwards, forming a stream of millikelvin-temperature atoms. These atoms are subsequently captured by a second-stage cooling system in a separate chamber: a five-beam, red BB MOT operating on the narrow $\rm{}^{1}S_{0}\rightarrow\,^{3}P_{1}$ transition [see Fig. 1(a)]. The red MOT quadrupole field is generated by a coil pair of approximately 20 cm diameter, arranged in an anti-Helmholtz configuration with vertical axis. The MOT laser beams are frequency modulated by a single-pass AOM, which is driven by a triangle wave modulated RF signal. The RF signal is created by a DDS (Analog Devices AD9852) and has a modulation bandwidth of $f_{\text{bandwidth}}=1.5\,$ MHz and a modulation frequency of $f_{\text{mod}}=40\,$kHz (see Fig. 2, inset). Further details on the apparatus and methodology can be found in Ref. [1].
@@
When using the BB MOT position as a frequency reference, we lock the MOT laser source to the frequency comb (A.1). The comb repetition rate is synchronized to a tunable, low-noise 800 MHz oscillator (TOPTICA RF TUNE),which serves as a short-term reference, while long-term frequency stability is transferred via the MOT position to the tunable RF oscillator (A.2). To enable continuous feedback, we switch to non-destructive fluorescence imaging. The MOT position is extracted every 50 ms via live fitting of the fluorescence image. Using the known frequency-to-position conversion factor $\Delta f/\Delta z$, a Python-based proportional-integral (PI) loop continuously feeds back to the tunable RF oscillator. This stabilization scheme, utilizing the comb, effectively transfers frequency stability to both the RF and optical domains simultaneously. We monitor both using a K+K counter (FXE Phase + Frequency Meter); the RF is mixed down and measured directly, while the optical stability is measured via a beat note between the DL Pro (locked to the MOT position) and our conventional laser system. The K+K counter is synchronized to a stable 10 MHz RF reference from the Dutch Metrology Institute (VSL), which has an fractional instability of approximately $10^{-14}$ after 100 seconds of averaging [8].
@@
Figure 2: Frequency sensitivity of the broadband ($B$) and single-frequency (SF) MOTs, demonstrating MOT laser linewidth independence. Despite differing forces, both MOTs exhibit similar sensitivities within error bars. Triangles: vertical magnetic field gradient $0.47\,\mathrm{G/cm}$, sensitivities $98.19(30)\,\mathrm{Hz/\mu m}$ ($B$, blue) and $98.76(50)\,\mathrm{Hz/\mu m}$ (SF, red). Crosses: $0.37\,\mathrm{G/cm}$, with $77.91(30)\,\mathrm{Hz/\mu m}$ (light blue, $B$) and $77.21(40)\,\mathrm{Hz/\mu m}$ (light red, SF). Stars: $0.27\,\mathrm{G/cm}$, with $57.37(24)\,\mathrm{Hz/\mu m}$ (light orange, $B$) and $57.25(40)\,\mathrm{Hz/\mu m}$ (turquoise, SF). An offset magnetic field shifts the intersection point of the six curves to $\delta_{\text{start}}=1.05\,\mathrm{kHz}$. Bottom inset: MOT AOM RF spectra showing $B$ modulation (blue) and SF (red) tone. Top inset: high-resolution BB MOT position scan near the intersection point, with frequency resolution below $250$\,\mathrm{Hz}$ at $0.31\,\mathrm{G/cm}$.
@@
Response Time. To evaluate the achievable feedback bandwidth for continuous measurement of the MOT position, it is essential to characterize the system’s response time i.e., the time required for the MOT to reach a new equilibrium after a change in MOT laser frequency. Figure 3 shows the BB MOT position as a function of time following frequency jumps of 10 kHz (red disks), 50 kHz (light red triangles), 150 kHz (orange crosses), and 300 kHz (light orange stars), using the same parameters we later employ in the long-term stability measurement. The MOT’s motion can be described by an overdamped harmonic oscillator model, with the position evolving as $z(t)=z_{\text{start}}+\left(1-e^{-t/\tau}\right)(z_{\text{end}}-z_{\text{start}})$, where $z_{\text{start}}$ and $z_{\text{end}}$ denote the initial and final MOT positions, and $\tau$ is the characteristic response time. As shown in the inset of Fig. 3, $\tau$ increases with the magnitude of the frequency jump. We model this dependence empirically as $\tau(f)=A\sqrt{f}+\tau_{\text{offset}}$. For a frequency change of less then 100 kHz, the system reaches equilibrium in about 25 ms, corresponding to a maximum feedback bandwidth of roughly 40 Hz.
@@
Continuous MOT position spectroscopy. Finally, we implement continuous MOT position spectroscopy using fluorescence imaging of the BB MOT (10 ms exposure time, 20 Hz repetition rate). The red MOT laser is stabilized to a frequency comb, and the feedback loop is closed by transferring the MOT position stability to the tunable RF oscillator, which provides the frequency reference for the comb. We compare the overlapping Allan deviation (ADEV with $100\,\mathrm{ms}$ gate time) for a total measurement time of $T=4600\,\mathrm{s}$. Error bars indicate $1\sigma$ confidence intervals, calculated using a chi-squared distribution with conservative $T/\tau-1$ degrees of freedom where $\tau$ is the cluster time [26]. Comparing the free-running RF oscillator to the MOT position stabilized case, we observe an improvement of more than two orders of magnitude after 400 s of averaging (see Fig. 4).The system reaches a minimum fractional frequency instability of $4.4\times 10^{-13}$ at 400 s, mainly limited by magnetic field fluctuations. Notably, we see that the MOT position stabilized tunable RF oscillator can outperform the beat signal between the MOT laser (locked via the comb to the MOT position) and the conventional laser system in the range of 100 to 1000 s. This points to a limitation of the conventional laser system due to the hot vapor spectroscopy. These results demonstrate that MOT position spectroscopy can exceed the long-term frequency stability of hot vapor modulation transfer spectroscopy, even when relying solely on passive stability for magnetic field and atom number. With the addition of active stabilization, the system could approach the performance limit set by the in-loop error signal, calculated by converting MOT positions (recorded with feedback on) into frequency space via the frequency-to-position conversion factor. Therefore, the Allan deviation is expected to follow $\sigma(\tau)=(6.71\pm 0.48)\times 10^{-12}/\tau^{0.84\pm 0.03},$ potentially reaching the high $10^{-14}$ range after just over 100 s of averaging (see Appendix A.5), where we also expect the noise floor to be limited by magnetic field fluctuations.
@@
Summary and outlook. We have demonstrated, to the best of our knowledge, the first continuous spectroscopy based on the position of a narrow-line strontium MOT, achieving a frequency sensitivity 30 times narrower than the natural linewidth of the transition. This method provides long-term frequency stability surpassing that of hot vapor modulation transfer spectroscopy, while maintaining a locking range and effective laser linewidth two orders of magnitude broader than the transition’s natural linewidth. If we actively stabilize the MOT position by feeding back to the repetition rate of a dispersion-optimized frequency comb [1], we reach a minimum fractional instability of less than $7\times 10^{-13}$ after 400 s of averaging in both the RF and optical domains, primarily limited by magnetic field fluctuations. With magnetic shielding [24] or active stabilization of stray environmental fields [25], it is possible to reduce the uncertainty even closer to the limit set by the in-loop error signal. Further improvement of the short-term stability could be achieved by synchronizing the tunable RF oscillator to a lower noise quartz oscillator [27] or a four-quadrant photodiode for higher-bandwidth MOT position measurements. On the other hand, one could also leverage the magnetic field sensitivity of this spectroscopy method and use such a system as a magnetometer in the direction of gravity [28].

### v2

@@ Abstract
We demonstrate a continuous spectroscopy technique with frequency sensitivity well below the natural transition linewidth, while maintaining a locking range hundreds of times larger. The method exploits the position dependence of a continuous, broadband magneto-optical trap operating on the 7.5 kHz-wide intercombination line of strontium. We show that the frequency sensitivity is fundamentally insensitive to the effective MOT laser linewidth. By applying active feedback on the MOT position to a dispersion-optimized frequency comb, which serves as the reference for stabilizing the MOT laser [4], we achieve a frequency instability below $4.4\times 10^{-13}$ after 400 s of averaging in both the optical and radio-frequency domains, surpassing the performance of conventional hot-vapor modulation transfer spectroscopy. Our method is a broadly applicable alternative route to frequency references in the high $10^{-14}$ range around 100 s.
@@
Continuously operating atomic clocks, such as hydrogen masers in the RF domain [28, 17], and systems employing modulation transfer spectroscopy in the optical domain [22, 24, 18, 7], have demonstrated frequency stability on the order of $\sim 10^{-14}$ after 100 s of averaging [22, 24, 18, 7, 6]. Recent advances have enabled the development of compact, field-deployable optical atomic clocks, including systems based on tetrahedral and pyramidal magneto-optical traps (MOTs) [26, 34, 2], microcell-based clocks [39, 16], and vapor-based clocks [21, 12, 9]. While modulation transfer spectroscopy remains widely used in these systems, it is limited by short-term probe laser instability and a narrow frequency locking range, both typically on the order of the transition linewidth [29, 42]. These limitations are especially pronounced for narrow-line optical transitions, such as those found in the alkaline-earth and alkaline-earth-like atoms. Additionally, modulation transfer spectroscopy does not provide direct access to the RF domain.
@@
In this Letter, we present a simple, continuous spectroscopy technique that exploits the dependence of the MOT position along the vertical axis (gravity) on the MOT laser frequency in a five-beam configuration. We demonstrate that this method improves both the frequency locking range and tolerance to short-term probe laser instability by two orders of magnitude compared to the natural linewidth, while simultaneously achieving a spectral resolution more than an order of magnitude below the 7.5 kHz natural linewidth of the $\mathrm{{}^{88}Sr}$ strontium intercombination transition. By applying active feedback based on the MOT position to a dispersion-optimized frequency comb, used for stabilizing the MOT laser, we extract both an 800 MHz RF reference and an optical reference at 689 nm, reducing frequency instabilities below $4.4\times 10^{-13}$ after 400 s of averaging. Finally, we demonstrate that this technique can outperform modulation transfer spectroscopy with hot atomic vapors in terms of averaging times of $\sim 100~\mathrm{s}$ on the same transition.
@@
Theory. To model the vertical dynamics of the five-beam MOT, where in comparison to a standard MOT the sixth beam propagating in gravity direction is removed [4], we consider the balance between gravitational force $mg$ and optical force along the vertical axis for a single atom with mass $m$ [23]. The resulting steady-state condition determines the MOT position as a function of laser detuning and magnetic field gradient [19, 33]. For strontium, using the $\rm{}^{1}S_{0}\rightarrow\,^{3}P_{1}$ intercombination line at 689 nm with a narrow linewidth of $\gamma_{0}=2\pi\cdot 7.5\,\mathrm{kHz}$, power broadening alone is insufficient to achieve large capture velocities. Instead, frequency modulation is used to extend the effective capture range by more than an order of magnitude. While this “broadband (BB) MOT” technique increases the capture velocities, it cannot be modeled simply as a power-broadened single-frequency (SF) light force. In our implementation, the RF drive of the AOM is triangularly modulated between $\delta_{\text{start}}$ and $\delta_{\text{stop}}=\delta_{\text{start}}-2\pi f_{\text{bandwidth}}$, where both $\delta$ values denote detunings from resonance, and $f_{\text{bandwidth}}$ is the total modulation range [1, 4]. In the low-saturation regime (see Appendix A.4), the total force in the vertical direction is:
@@
where $n=\mathrm{round}\!\left(f_{\text{bandwidth}}/f_{\text{mod}}\right)$ is the total number of frequency components considered, $f_{\text{mod}}$ is the modulation frequency, $\sum_{i=0}^{n}\alpha_{i}=1$ are the relative intensities of the $i^{\text{th}}$ frequency component, $I_{\text{total}}=2P/(\pi\omega_{0}^{2})$ is the total intensity of the MOT beam (defined by the beam power $P$ and gaussian waist size $\omega_{0}$) and $I_{0}$ the saturation intensity. For $m_{J}=\pm 1$ of this transition, the Zeeman shift is $\frac{\mu_{B}g_{J}}{h}=2101~\text{kHz/G}$ [1], where $\mu_{B}$ is the Bohr magneton and $g_{J}$ is the Landé g-factor. In equilibrium, $F_{\text{total}}=0$, and the MOT position is directly related to the Zeeman shift. Since we operate close to the center of a quadrupole magnetic field with a linear gradient $B^{\prime}$, a change $\delta_{\text{start}}\rightarrow\delta_{\text{start}}+2\pi\Delta f$ results in a corresponding vertical displacement $\Delta z$, where the frequency-to-position conversion factor $\Delta f/\Delta z=\frac{\mu_{B}g_{J}B^{\prime}}{\hbar}$ is independent of natural linewidth $\gamma_{0}$ and power broadening.
@@
Figure 1: Experimental schematic: (a) Atoms (green disks) are continuously loaded from an upper 2D blue MOT ((b), operating on the $\rm{}^{1}S_{0}\rightarrow^{1}P_{1}$ transition at 461 nm) into a five-beam broadband $\mathrm{{}^{88}Sr}$ red MOT ((b), operating on the narrow $\rm{}^{1}S_{0}\rightarrow^{3}P_{1}$ transition at 689 nm). The red MOT is monitored non-destructively via fluorescence imaging, and the center position along the vertical (gravity) axis is extracted every 50 ms. The red MOT laser is locked to an optical frequency comb, which provides short-term frequency stability, while long-term stability is achieved via the MOT position where the vertical displacement $\delta z$ serves as an error signal. Feedback is applied through a tunable RF oscillator used to synchronize the comb, which in turn transfers its stability to both the RF and optical domains.
@@
Experimental Setup. We employ our previously developed architecture for generating a narrow-line $\mathrm{{}^{88}Sr}$ MOT using a dispersion-optimized frequency comb-locked laser system, without the need for a free space high-finesse optical cavity [4]. In contrast to typical experiments that execute cooling stages time-sequentially, we execute them in a spatial sequence. Atoms effuse from an oven, are Zeeman slowed and then cooled in a blue 2D MOT, out of which they continuously fall downwards, forming a stream of millikelvin-temperature atoms. These atoms are subsequently captured by a second-stage cooling system in a separate chamber: a five-beam, red BB MOT operating on the narrow $\rm{}^{1}S_{0}\rightarrow\,^{3}P_{1}$ transition [see Fig. 1(a)]. The red MOT quadrupole field is generated by a coil pair of approximately 20 cm diameter, arranged in an anti-Helmholtz configuration with vertical axis. The MOT laser beams are frequency modulated by a single-pass AOM, which is driven by a modulated RF signal. This signal is generated by a DDS (Analog Devices AD9852) with a triangular frequency modulation with a total modulation range of $f_{\text{bandwidth}}=1.5~\mathrm{MHz}$ and modulation frequency $f_{\text{mod}}=40~\mathrm{kHz}$ (see Fig. 2, inset). Further details on the apparatus and methodology can be found in Ref. [4].
@@
When using the BB MOT position as a frequency reference, we lock the MOT laser source to the frequency comb (A.1). The comb repetition rate is synchronized to a tunable, low-noise 800 MHz oscillator (TOPTICA RF TUNE),which serves as a short-term reference, while long-term frequency stability is transferred via the MOT position to the tunable RF oscillator (A.2). To enable continuous feedback, we switch to non-destructive fluorescence imaging. The MOT position is extracted every 50 ms via live fitting of the fluorescence image. Using the known frequency-to-position conversion factor $\Delta f/\Delta z$, a Python-based proportional-integral (PI) loop continuously feeds back to the tunable RF oscillator. This stabilization scheme, utilizing the comb, effectively transfers frequency stability to both the RF and optical domains simultaneously. We monitor both using a K+K counter (FXE Phase + Frequency Meter); the RF is mixed down and measured directly, while the optical stability is measured via a beat note between the DL Pro (locked to the MOT position) and our conventional laser system. The K+K counter is synchronized to a stable 10 MHz RF reference from the Dutch Metrology Institute (VSL), which has an fractional instability of approximately $10^{-14}$ after 100 seconds of averaging [6, 38, 37, 4].
@@
Figure 2: Frequency sensitivity of the broadband ($B$) and single-frequency (SF) MOTs, demonstrating MOT laser linewidth independence. Despite differing forces, both MOTs exhibit similar sensitivities within error bars. Triangles: vertical magnetic field gradient $0.47\,\mathrm{G/cm}$, sensitivities $98.19(30)\,\mathrm{Hz/\mu m}$ ($B$, blue) and $98.76(50)\,\mathrm{Hz/\mu m}$ (SF, red). Crosses: $0.37\,\mathrm{G/cm}$, with $77.91(30)\,\mathrm{Hz/\mu m}$ (light blue, $B$) and $77.21(40)\,\mathrm{Hz/\mu m}$ (light red, SF). Stars: $0.27\,\mathrm{G/cm}$, with $57.37(24)\,\mathrm{Hz/\mu m}$ (light orange, $B$) and $57.25(40)\,\mathrm{Hz/\mu m}$ (turquoise, SF). An offset magnetic field shifts the intersection point of the six curves to $\delta_{\text{start}}=1.05\,\mathrm{kHz}$. Bottom inset: MOT AOM RF spectra showing $B$ modulation (blue) and SF (red) tone.
Top inset: high-resolution BB MOT position scan near the intersection point, with frequency resolution below $250$\,\mathrm{Hz}$ at $0.31\,\mathrm{G/cm}$.
@@
Response Time. To evaluate the achievable feedback bandwidth for continuous measurement of the MOT position, it is essential to characterize the system’s response time i.e., the time required for the MOT to reach a new equilibrium after a change in MOT laser frequency. Figure 3 shows the BB MOT position as a function of time following frequency jumps of 10 kHz (red disks), 50 kHz (light red triangles), 150 kHz (orange crosses), and 300 kHz (light orange stars), using the same parameters we later employ in the long-term stability measurement (averaging times of $\sim 100~\mathrm{s}$). The MOT’s motion can be described by an overdamped harmonic oscillator model, with the position evolving as $z(t)=z_{\text{start}}+\left(1-e^{-t/\tau}\right)(z_{\text{end}}-z_{\text{start}})$, where $z_{\text{start}}$ and $z_{\text{end}}$ denote the initial and final MOT positions, and $\tau$ is the characteristic response time. As shown in the inset of Fig. 3, $\tau$ increases with the magnitude of the frequency jump. We model this dependence empirically as $\tau(f)=A\sqrt{f}+\tau_{\text{offset}}$. For a frequency change of less then 100 kHz, the system reaches equilibrium in about 25 ms, corresponding to a maximum feedback bandwidth of roughly 40 Hz.
@@
Continuous MOT position spectroscopy. Finally, we implement continuous MOT position spectroscopy using fluorescence imaging of the BB MOT (10 ms exposure time, 20 Hz repetition rate). The red MOT laser is stabilized to a frequency comb, and the feedback loop is closed by transferring the MOT position stability to the tunable RF oscillator, which provides the frequency reference for the comb. We compare the overlapping Allan deviation (ADEV with $100\,\mathrm{ms}$ gate time) for a total measurement time of $T=4600\,\mathrm{s}$. Error bars indicate $1\sigma$ confidence intervals, calculated using a chi-squared distribution with conservative $T/\tau-1$ degrees of freedom where $\tau$ is the cluster time [32]. Comparing the free-running RF oscillator to the MOT position stabilized case, we observe an improvement of almost two orders of magnitude after 400 s of averaging (see Fig. 4).The system reaches a minimum fractional frequency instability of $4.4\times 10^{-13}$ at 400 s, mainly limited by magnetic field fluctuations. Notably, we see that the MOT position stabilized tunable RF oscillator can outperform the beat signal between the MOT laser (locked via the comb to the MOT position) and the conventional laser system in the range of 100 to 1000 s. This points to a limitation of the conventional laser system due to the hot vapor spectroscopy. These results demonstrate that MOT position spectroscopy can exceed the frequency stability of hot vapor modulation transfer spectroscopy for narrow intercombination lines in alkaline-earth(-like) atoms for averaging times of $\sim 100~\mathrm{s}$ with the potential for even longer time scales, even when relying solely on passive stability for magnetic field and atom number. With the addition of active stabilization, the system could approach the performance limit set by the in-loop error signal, calculated by converting MOT positions (recorded with feedback on) into frequency space via the frequency-to-position conversion factor. Therefore, the Allan deviation is expected to follow $\sigma(\tau)=(6.71\pm 0.48)\times 10^{-12}/\tau^{0.84\pm 0.03},$ potentially reaching the high $10^{-14}$ range after just over 100 s of averaging (see Appendix A.5), where we also expect the noise floor to be limited by magnetic field fluctuations.
@@
Summary and outlook. We have demonstrated, to the best of our knowledge, the first continuous spectroscopy based on the position of a narrow-line strontium MOT, achieving a frequency sensitivity 30 times narrower than the natural linewidth of the transition. This method provides frequency stability surpassing that of hot vapor modulation transfer spectroscopy for averaging times of $\sim 100~\mathrm{s}$ with the potential for even longer time scales, while maintaining a locking range and effective laser linewidth two orders of magnitude broader than the transition’s natural linewidth. If we actively stabilize the MOT position by feeding back to the repetition rate of a dispersion-optimized frequency comb [4], we reach a minimum fractional instability of less than $7\times 10^{-13}$ after 400 s of averaging in both the RF and optical domains, primarily limited by magnetic field fluctuations. With magnetic shielding [35] or active stabilization of stray environmental fields [3], it is possible to reduce the uncertainty even closer to the limit set by the in-loop error signal. Further improvement of the short-term stability could be achieved by synchronizing the tunable RF oscillator to a lower noise quartz oscillator [31] or a four-quadrant photodiode for higher-bandwidth MOT position measurements. On the other hand, one could also leverage the magnetic field sensitivity of this spectroscopy method and use such a system as a magnetometer in the direction of gravity [40].